# Causal LLM Routing: End-to-End Regret Minimization from Observational Data

### v1

@@ 5 Conclusion
We propose a causal end-to-end framework for routing queries to LLMs under observational data. Our approach introduces a regret-minimizing objective grounded in counterfactual estimation, enabling principled policy learning that accounts for treatment selection bias without requiring full-feedback data. Unlike prior approaches that rely on decoupled prediction of accuracy and cost, where errors can compound, our method directly optimizes the decision objective. To support heterogeneous user preferences, we develop an interval-conditioned routing architecture that generalizes across a continuum of cost-sensitivity parameters. Theoretical analysis provides guarantees on interpolation sufficiency and regret bounds, while empirical evaluations on public routing benchmarks demonstrate that our methods consistently outperform strong baselines, including recent routing algorithms, across multiple embedding models. Future work includes extending the framework to accommodate additional user-defined metrics or hard constraints that cannot be readily incorporated as soft penalties in the objective. Another promising direction is exploring online or adaptive routing in dynamic environments, as well as extending causal regret minimization to multi-turn settings.
@@ Appendix A Proofs of Propositions
Under Assumption 1, the regret approximation using the doubly robust estimator is upper bounded by
$\text{Regret}(f)\leq L\cdot\frac{1}{n}\sum_{i=1}^{n}d(t_{i}^{*},f(x_{i})),$ (13)
where $t_{i}^{*}:=\arg\max_{t\in\mathcal{T}}\hat{Y}_{x_{i}}(t)$.
@@
We have that $Regret(f)\approx\frac{1}{n}\big{(}\sum_{i=1}^{n}\hat{Y}_{x_{i}}(t^{*}_{i})-\hat{Y}_{x_{i}}(f(x_{i}))\big{)}$. Thus, by the $L$-Lipschitz continuity of $\hat{Y}_{x_{i}}(\cdot)$, we have that $\frac{1}{n}\sum_{i=1}^{n}\big{(}\hat{Y}_{x_{i}}(t_{i}^{*})-\hat{Y}_{x_{i}}(f(x_{i})\big{)}\leq L\cdot\frac{1}{n}\sum_{i=1}^{n}d(t_{i}^{*},f(x_{i}))$, which yields the claimed upper bound. ∎
@@ B.2 Additional Tables
[…]
We also report the AUC (Area Under the Curve) of the accuracy–cost trade-off curve for each method. This is computed using the sklearn.metrics.auc function Buitinck et al. [2013]. AUC provides a single scalar summary of performance that captures how well a model balances accuracy and computational cost. A higher AUC indicates a more favorable overall trade-off across budgets, making it a robust evaluation metric for comparing routing strategies.
@@
[…]
Appendix C Experimental Details
All experiments were implemented in Python 3.8.12 Van Rossum and Drake [2009], using PyTorch 2.4.1+cu121 Paszke et al. [2019] and Scikit-learn Buitinck et al. [2013]. Experiments were conducted on an internal compute cluster equipped with an Intel(R) Xeon(R) Platinum 8260 CPU @ 2.40GHz, 512 GB of RAM, and two NVIDIA V100 GPUs with 16 GB memory each.
@@ Appendix C Experimental Details
RouterBench [Hu et al., 2024] is a standardized benchmark for LLM routing, comprising 35,712 prompt-response pairs collected from 11 LLMs. The prompts span eight different evaluation benchmarks covering reasoning, factual knowledge, dialogue, mathematics, and code generation. Each prompt is annotated with model accuracy and execution cost, enabling response-based decision-making. To maintain consistency in evaluation, we adopt the same split strategy for RouterBench, applied deterministically at the prompt level to ensure reproducibility. SPROUT [Somerstep et al., 2025] is a more recent and larger benchmark for cost-aware routing, consisting of 44,241 prompts and responses from 13 state-of-the-art language models. The prompts are drawn from six diverse benchmarks, including GPQA [Rein et al., 2024], MuSR [Sprague et al., 2023], MMLU-Pro [Wang et al., 2024], MATH [Hendrycks et al., 2021], OpenHermes [Teknium, 2023], and RAGBench [Friel et al., 2024]. SPROUT includes a predefined train/validation/test split, using 80% of the data for training and splitting the remaining 20% equally between validation and test sets.

### v2

@@ 5 Conclusion
We propose a causal end-to-end framework for routing queries to LLMs under observational data. Our approach introduces a regret-minimizing objective grounded in counterfactual estimation, enabling principled policy learning that accounts for treatment selection bias without requiring full-feedback data. Unlike prior approaches that rely on decoupled prediction of accuracy and cost, where errors can compound, our method directly optimizes the decision objective. To support heterogeneous user preferences, we develop an interval-conditioned routing architecture that generalizes across a continuum of cost-sensitivity parameters. Theoretical analysis provides guarantees on interpolation sufficiency and regret bounds, while empirical evaluations on public routing benchmarks demonstrate that our methods consistently outperform strong baselines, including recent routing algorithms, across multiple embedding models. Future work includes extending the framework to accommodate additional user-defined metrics or hard constraints that cannot be readily incorporated as soft penalties in the objective. Another promising direction is to explore online or adaptive routing in dynamic environments, as well as extending causal regret minimization to multi-turn settings.
app
@@ Appendix A Proofs of Propositions
Suppose the estimated utility function $\hat{Y}_{x}:\mathcal{T}\to\mathbb{R}$ is $L$-Lipschitz continuous over the probability simplex with respect to the $\ell_{1}$ norm, as in Definition 1. Then, for a policy $f:\mathcal{X}\to\Delta^{|\mathcal{T}|}$ that outputs a distribution $f(x)$ over $\mathcal{T}$, the regret can be upper bounded by:
$\text{Regret}(f)\leq L\cdot\frac{1}{n}\sum_{i=1}^{n}\sqrt{2\cdot\text{CE}(t_{i}^{*},f(x_{i}))},$ (15)
where $t_{i}^{*}:=\arg\max_{t\in\mathcal{T}}\hat{Y}_{x_{i}}(t)$ is the optimal treatment for input $x_{i}$, and $\text{CE}(t_{i}^{*},f(x_{i})):=-\log f(x_{i})_{t_{i}^{*}}$ denotes the cross-entropy loss.
@@
Let $e_{t_{i}^{*}}\in\Delta^{|\mathcal{T}|}$ denote the one-hot distribution over the optimal treatment $t_{i}^{*}$. By the definition of regret:
$\text{Regret}(f)$ $=\frac{1}{n}\sum_{i=1}^{n}\left[\hat{Y}_{x_{i}}(e_{t_{i}^{*}})-\hat{Y}_{x_{i}}(f(x_{i}))\right].$ (16)
Using the $L$-Lipschitz continuity of $\hat{Y}_{x_{i}}(\cdot)$ under the $\ell_{1}$ norm:
$\left|\hat{Y}_{x_{i}}(e_{t_{i}^{*}})-\hat{Y}_{x_{i}}(f(x_{i}))\right|$ $\leq L\cdot\|e_{t_{i}^{*}}-f(x_{i})\|_{1}.$ (17)
Applying Pinsker’s inequality:
$\|e_{t_{i}^{*}}-f(x_{i})\|_{1}$ $\leq\sqrt{2\cdot\text{KL}(e_{t_{i}^{*}}\,\|\,f(x_{i}))}=\sqrt{2\cdot\text{CE}(t_{i}^{*},f(x_{i}))},$ (18)
where the equality follows because KL divergence from a one-hot distribution to a probability vector reduces to cross-entropy. Combining the above:
$\text{Regret}(f)$ $\leq L\cdot\frac{1}{n}\sum_{i=1}^{n}\sqrt{2\cdot\text{CE}(t_{i}^{*},f(x_{i}))},$ (19)
which completes the proof. ∎
@@ B.2 Additional Tables
RM-Interval 84.60±0.89 76.54±4.06 71.45±6.01 69.99±0.62 67.18±0.51 64.73±0.97 63.04±1.35 61.16±1.84 59.04±1.05 57.48±2.90 57.12±0.21
We also report the AUC (Area Under the Curve) of the accuracy–cost trade-off curve for each method. This is computed using the sklearn.metrics.auc function sklearn_api. AUC provides a single scalar summary of performance that captures how well a model balances accuracy and computational cost. A higher AUC indicates a more favorable overall trade-off across budgets, making it a robust evaluation metric for comparing routing strategies.
@@
RM-Interval $0.2285$ $0.2196$ $0.3320$ $0.2464$
Appendix C Contribution of Each Component
The results highlight the impact of causal bias correction, end-to-end training, and regret-focused objectives as follows:
Causal Inference (Bias Correction):
Our Baseline corresponds to CARROT without any causal correction - that is, a decoupled predictor trained directly on observational data without accounting for treatment bias. It consistently performs the worst across all settings. In contrast, incorporating causal techniques for bias correction yields substantial gains: both R&C and Causal-CARROT, which integrate causal adjustments into routing, achieve +10–15% routing accuracy at comparable cost, demonstrating that accounting for selection bias is critical for effective routing. These results validate that bias-aware routing significantly improves utility over naive predictors trained on biased data.
End-to-End Learning vs. Two-Stage:
Among bias-corrected approaches, our end-to-end regret-minimizing methods (RM) consistently outperform the two-stage methods (CF-Regression, Causal-CARROT, R&C). The performance gap (+1–3% routing accuracy at comparable cost) demonstrates the benefit of an integrated end-to-end approach: by directly optimizing the decision-quality objective (regret) rather than optimizing intermediate predictions, our method achieves superior and more stable results.
Regret Minimization Objective:
Even compared to other causal learners, our specific training objective provides an edge. For instance, RM-Softmax (differentiable surrogate) slightly outperforms RM-Classification (upper-bound surrogate) in most cases, with lower variance (+0.5–1% routing accuracy at comparable cost). This highlights the advantage of our softmax-weighted regret objective and its alignment with the true decision loss. Moreover, both RM methods outperform Causal-CARROT and CF-Regression, underscoring that minimizing expected regret is more effective than surrogate approaches that focus only on accuracy/cost prediction.
Appendix D Experimental Details
All experiments were implemented in Python 3.8.12 python, using PyTorch 2.4.1+cu121 NEURIPS2019_9015 and Scikit-learn sklearn_api. Experiments were conducted on an internal compute cluster equipped with an Intel(R) Xeon(R) Platinum 8260 CPU @ 2.40GHz, 512 GB of RAM, and two NVIDIA V100 GPUs with 16 GB memory each.
@@ Appendix D Experimental Details
RouterBench [hu2024routerbench] is a standardized benchmark for LLM routing, comprising 35,712 prompt-response pairs collected from 11 LLMs. The prompts span eight different evaluation benchmarks covering reasoning, factual knowledge, dialogue, mathematics, and code generation. Each prompt is annotated with model accuracy and execution cost, enabling response-based decision-making. To maintain consistency in evaluation, we adopt the same split strategy for RouterBench, applied deterministically at the prompt level to ensure reproducibility. SPROUT [somerstep2025carrot] is a more recent and larger benchmark for cost-aware routing, consisting of 44,241 prompts and responses from 13 state-of-the-art language models. The prompts are drawn from six diverse benchmarks, including GPQA [rein2024gpqa], MuSR [sprague2023musr], MMLU-Pro [wang2024mmlu], MATH [hendrycks2021measuring], OpenHermes [teknium2023openhermes], and RAGBench [friel2407ragbench]. SPROUT includes a predefined train/validation/test split, using 80% of the data for training and splitting the remaining 20% equally between validation and test sets.
@@
Inference latency and computational efficiency
Latency is critical in real-time applications. Our method is designed to minimize this by using lightweight routing networks (2-layer MLP with 200 neurons per hidden layer) and precomputed light-weight embeddings (Llama-3.2-1B and BERT). For instance, with Llama-3.2-1B embeddings and an MLP-based router on the RoutherBench dataset, the end-to-end routing latency is under 2.5 ms on a single A100 GPU for a batch size of almost 25,000. To contextualize this, recent benchmarks show that on 8× A100s, LLaMA‑2 or LLaMA‑3 70B can generate a  100-token paragraph in 1.5 to 2.5 seconds under realistic conditions using optimized inference stacks. Even when using fewer GPUs or smaller models, generation latency typically remains one to two orders of magnitude higher than our routing time. In many cases (e.g., longer outputs or cold starts), the difference can be significantly larger. Thus, the routing overhead is negligible relative to the cost of LLM generation and does not meaningfully impact user experience. Additionally, the interval-conditioned architecture reduces deployment complexity by avoiding the need to train or store separate models for different user preferences.